THE SUBMILLIMETER-WAVE SPECTRUM OF TRANS- AND CIS-CROTONONITRILE (CH_{3}CH=CHCN)

\underline{BRIAN J. DROUIN} AND JOHN C. PEARSON, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109-8099.

The ground state rotational spectrum and ground state rotational torsional spectrum of trans- and cis-crotononitrile are further characterized in the 270-480 GHz frequency range. The A-E splittings due to methyl internal rotation in trans-crotononitrile^{a} are not observed in the sub-millimeter spectrum which has been assigned and fit to a semi-rigid rotor Hamiltonian in agreement with previous work^{b}. For cis-crotononitrile over 380 transitions are now assigned to this conformation, twice the number previously reported. Internal rotation splittings are observed throughout the^{a}R−branches and have been assigned for J > 30. The A-E spectrum fits reasonably well in an internal axis system, but strong correlation between D_{ab} and the rotational constants indicates that a rotated internal axis system may suit the data better.

aM. Suzuki and K. Kozima, J. Mol. Spec. 33, 407-413 (1970); S. L. Hsu and W. H. Flygare, J. Mol. Spec. 37, 92-99 (1971).

bA.G. Lesarri, J. Cosleou, X. Li, G. Wlodarczak and J. Demaison, J. Mol. Spec. 172, 520-535 (1995).